OpenStudy (anonymous):

Rearrange this formula to make M the subject: t = 2M - v

OpenStudy (anonymous):

M = (t+v)/2

OpenStudy (anonymous):

t=2M-v t+v=2M (t+v)/2=M

OpenStudy (anonymous):

so it's (t + v) divided by 2?